OpenStudy (anonymous):

write a equation of the line containing the given point and parallel to the given line express your answers in the form y+mx+b the equation of the line is y=? (7,9);x+9y=5

OpenStudy (tad1):

parallel lines have the same slope. So the slope of your new line will be the same as the first. x + 9y = 5 9y = -x + 5 y = (-1/9)x + 5/9 so m = -1/9 then use point-slope form y-9 = -(1/9)(x -7)